How are you supposed to substitute a equation if it one of them are not in y=mx+b form?
Example: x-y=1 2x+y=8
Answer by rapaljer(4671) 
 You'll have to solve for one of the variables in terms of the other one in one of the equations. In the example you gave, it would be pretty easy to solve for x in the first equation by adding +y to each side of the equation, so
x-y = 1, so x = 1+y
Now, substitute x = 1+y into the second equation as is:
2x + y = 8
2(_______) + y = 8
2(1+y) + y = 8
2 + 2y + y = 8
2+3y = 8
Subtract 2 from each side:
3y = 6
y = 2
Now, substitute y = 2, back into the equation where you solved for x: x= 1+y
x=1+y
x= 1+2
x= 3
Therefore the point (3,2) is a solution.
Check by substituting both numbers into the second equation:
2x+y = 8
2(3) + 2 = 8
6+2 = 8
It checks!!
R^2 at SCC
|
|
|